OpenStudy (anonymous):

Solve the proportion: 5/x = 8/15

OpenStudy (anonymous):

\[\frac{ 5 }{ x } = \frac{ 8 }{ 15 } \] \[\frac{ 5x }{ x } = \frac{ 8x }{ 15 }\] \[5 = \frac{ 8x }{ 15 }\] Swapping x to the left \[\frac{ 8x }{ 15 } = 5\] \[\frac{ 8x }{ 15 } \times \frac{ 15 }{ 8 } = 5 \times \frac{ 15 }{ 8 }\] \[x=\frac{ 75 }{ 8 }\]

OpenStudy (anonymous):